Additional Aspects of aqueous balances:

2. Determine the pH of a buffer solution consisting of 0.50M of CH3COOH and 0.50M of CH3COO- . Ka=1.8x10^-5
Answer #1

Concentration of CH3COOH = 0.50 M

Concentration of CH3COO- = 050 M

Acid dissociation constant Ka = 1.8 x 10^-5

pKa = - log Ka = - log (1.8 x 10^-5) = 4.74

From the Henderson-Hasselbalch equation

pH = pKa + log [(base)/(acid)]

pH = pKa + log [(CH3COO-)/(CH3COOH)]

pH = 4.74 + log (0.50/0.50)

pH =4.74 + 0

pH = 4.74